OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? x 1 2 3 4 y 2 7 12 17

Mehek (mehek14):

subtract y2-y1/x2-x1

Mehek (mehek14):

choose any two points

OpenStudy (anonymous):

1- 1/5x+b

Mehek (mehek14):

?

OpenStudy (anonymous):

m=y2-y1 x2-x1 right

Mehek (mehek14):

yes so choose 2 points

OpenStudy (anonymous):

m=2/7-1/2=1/5

Mehek (mehek14):

it would be the other way 7/2-2/1 = 5/1 because y has to be on top and x on the bottom

OpenStudy (anonymous):

O So What Do I Do Now

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

so your slope is 5 y = 5x + b now we have to find b substitute the value of any points

Mehek (mehek14):

can you do that?

OpenStudy (anonymous):

1-1/5=1/5- 1/5+b

Mehek (mehek14):

no not like that

Mehek (mehek14):

lets use 1,2 so since 2 is y and 1 is x, 2 = 1x + b now solve that

OpenStudy (anonymous):

1-1/2=1/2-1/2+b

Mehek (mehek14):

how do get that???

OpenStudy (anonymous):

i dont get what you have to do tbh

Mehek (mehek14):



Mehek (mehek14):

im using points 1,2 1 is x so substitute that for x and 2 is y so substitute that for y

OpenStudy (anonymous):

so thats the answer

Mehek (mehek14):

yes the equation

Mehek (mehek14):

do you get it?